2 Litres of mercury at 80 degree celsius is pushed into a plastic bucket containing 10 litres of mercury at temprature T. If the final temperature of the mixture is 30 degree celsius. What is the value of T?

Grade:12th Pass

1 Answers

suryakanth AskiitiansExpert-IITB
13 years ago

Dear vikas,

the heat lost by 80 degree mercury is taken by mercury at T degrees

using

M1Cp (80-30) = M2Cp (T-30)

density *volume *(50) = Density * volume (T-30)

2 *50 = 10 * (T-30)

=> 10 = T -30

=> T = 40 degrees

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Suryakanth –IITB